### v1

# A comprehensive theory for relativistic polaritonic chemistry: a way toward a four component ab initio treatment of molecular systems

Abstract
We present a new ab initio approach to study molecules containing heavy atoms strongly interacting with quantum fields in optical devices. The relativistic quantum electrodynamics (QED) theory has been rewritten with a formalism close to relativistic quantum chemistry. This general framework represents the ideal starting point to extend the main quantum chemistry methods to relativistic QED. The Polaritonic Dirac Hartree Fock (Pol-DHF) approach is the first method we propose based on this theory. Pol-DHF allows for the simulation of field induced effects on the ground and excited state properties of heavy transition metals molecular complexes. The method is able to include not only the effects of the photons but can be easily extended also to include explicit interactions with positrons. Application of Pol-DHF to three metal hydrides revealed the importance of including radiative QED corrections to the treatment in strong coupling conditions. Due to an accurate description of spin-orbit coupling, the method is able to reproduce polaritonic effects happening at the crossing between singlet and triplet potential energy surfaces.
I Introduction
The use of light as a new tool to control and manipulate non-invasively the properties of molecules and materials is opening, in recent years, a new field of research at the border between physics, chemistry and material science [1, 2, 3, 4, 5].
When matter strongly couples to photons, new hybrid states (polaritons), having partial light and partial matter character, are formed. The strong coupling condition is usually reached inside properly designed optical devices. The simplest example is the Fabry-Pérot cavity [6] (Fig 1), made of two highly reflective planar mirrors, that confine the photons in a small quantization volume and leading to a significant enhancement of the light-matter coupling.
The field properties and therefore the polaritons properties can be controlled changing the geometry and the materials of the cavity. Several demonstrations of polaritonic effects on different physical properties such as absorption spectra, photochemical reaction rates, and conductivity have already been observed in the experiments [2, 7, 3, 5, 8, 9, 10].
Probably the most famous demonstration was obtained, in the experiments performed by the Ebbesen’s group in Strasbourg. In particular, they demonstrated that strong coupling to molecular vibrations can be used to catalyze, slow down or even induce selectivity in chemical reactions [11, 5, 12]. These observations opened a new field that is now known as polaritonic chemistry [13].
In these experiments, the photonic states are usually coupled either to electronic or vibrational states of the molecular systems. However, the electromagnetic nature of the field allows also for modifications of their magnetic properties if a coupling to the spin states is exploited. In this way, a fine control of the magnetizability and aromatic properties of molecules [14], of spin qubits [15, 16] and of spin phases of materials [17, 18] can be obtained. However, reaching the strong coupling condition in this frequency range is impossible using a simple Fabry-Pérot cavity that, in this case, would require a spacing of centimeters between the mirrors. The problem can be circumvented using planar superconducting devices commonly used in Circuit-QED experiments [19] . Using similar devices Affronte’s group has been able to manipulate the spin properties of a Single-Molecule Magnet [16, 20]. Such an accomplishment opened large possibilities of application in spin qubits based quantum computation.
Despite the many improvements in the fabrication of more effective optical devices and the impressive accuracy reached [21, 22] by polaritonic chemistry experiments, many fundamental aspects still remain to be understood. In this context, theory represents a fundamental tool to gain insights on the underlying physics of these processes. In recent years many ab initio methods, able to treat electron-electron and electron-photon correlation at different level of accuracy have been developed [23, 24, 25, 26, 27, 28]. However, they have been mainly applied to investigate cavity induced effects on the electronic and vibrational degrees of freedom of molecular systems. Only very recently, Barlini et al. proposed the first Hartree-Fock based approach to study photon induced effects on the electronic and nuclear magnetic properties of a molecular system [14].
Spin and consequently magnetic properties are intrinsic relativistic features of molecules. An accurate investigation of these properties requires then inclusion of relativistic effects at different level of accuracy. This becomes particularly crucial when, like in Affronte’s experiments [16, 20] Single-Molecule Magnets containing lanthanides atoms need to be used to have sufficiently long living magnetizations.
Moreover, as already discussed in Refs [28, 29] for Van-der Waals interactions, the field has sometimes to power to enhance small effects that are usually neglected in absence of the photons. From these considerations rises the emergency of a consistent relativistic quantum electrodynamical ab initio theory able to include all the necessary effects. Formulating such a general approach is one of the goals of this paper. Similar intent was already presented by Ruggenthaler et al. [23] in a Density Functional Theory (DFT) framework, but only the non-relativistic limit of the method was actually turned into a usable quantum chemistry implementation. Very recently, Konecny et al. [30] proposed a relativistic response theory to investigate electronic excitations of relativistic molecules in optical cavities. Despite this last approach was able to capture interesting effects like cavity-induced singlet-triplet interactions, it is neglecting additional photon driven effects that might be crucial if magnetic properties want to be investigated. Moreover, this approach does not account for modifications induced by the field on the system’s ground state. This is a crucial aspect if we are interested to cavity driven effects on the core properties of relativistic atoms.
In this paper, we develop the first fully relativistic ab initio method to describe the ground state of molecular systems coupled to photons in optical cavities. We start the development from the most general Quantum Electrodynamical Lagrangian formulation to then propose a new and totally method independent approach to develop relativistic ab initio theories for integrating electron-photon systems. This formalism represents the starting point for the development of mean-field but also correlated methodologies. This strategy allows for a more controlled and conscious applications of approximations.
In the current state-of-the-art of relativistic quantum chemistry, the main QED corrections include electron-electron retardation effects (i.e. Breit), and additional radiative QED effects (i.e. vacuum polarization [31, 32], self-energy [33]) responsible for the Lamb-Shift [34]. These additional contributions are divergent and require regularization schemes (e.g. renormalization of the electron mass [35, 36]). Their inclusion in many-electrons systems is usually carried out using effective QED potentials.[37, 38, 39, 40, 41]. In a recent paper [42], it has been demonstrated that Breit contribution and these QED corrections effects on the Ionization Potential (IP) and Electron Affinities (EA) of gold are of the same order of magnitude (0.01-0.03 eV). In our case, instead, interactions with real cavity photons are comparable ( eV for the IP) [43] or (for heavy atoms) smaller than these radiative QED corrections. This clearly prevent us from neglecting a priori these effects. This is a crucial point that will be discussed in this paper and that motivates the proposed general reformulation of the relativistic polaritonic framework.
The paper is structured as follows: in section II a generic derivation of a relativistic QED Hamiltonian theory will be presented, starting from the standard QED Lagrangian. In this section, the choice of the gauge is discussed in detail and the different energy contributions are analyzed. At this stage, no approximations are used in order to set up a theory that could be the starting point for every quantum chemistry modelization. In subsection II.2.2 the coulomb gauge Hamiltonian, after application of the dipole approximation, is used to develop the first Relativistic Polaritonic HF (Pol-DHF) approach. In this context, large space has been dedicated to strategies used to deal with negative energy states and to improve convergence issues. In section III Pol-DHF has been applied to investigate electronic properties of small diatomic molecules containing heavy atoms. We end the paper with conclusions and perspectives.
II Theory
In this section, we follow the formal derivation of relativistic QED theory usually presented in physics text books[44, 45] to develop a Hamiltonian formalism that can be applied to formulate new ab initio methodologies for the simulation of polaritonic molecular systems. This approach allows to keep track of all the approximations used during the derivation, consequently facilitating future improvement strategies, as well as the overall comprehension of the method. The proposed methodology is then used to develop the first Hartree-Fock (HF) based approach for relativistic molecular systems strongly coupled to quantum fields. For convenience reasons, Gaussian units will be used through the whole derivation unless specified otherwise.
II.1 The Quantum Electrodynamics Lagrangian
We start our derivation from the definition of a Lagrangian describing at the same time the relativistic molecular system, the electromagnetic field and their interaction. In the following, all greek letters indices span the components of 4 indices vectors (from 0 to 3), whereas latin letters only span the spatial components of the vector (from 1 to 3).
To describe matter, we use the standard Lagrangian density for Dirac fields:
where is the speed of light and is the mass of the electron. The matrices are defined as:
with representing the Pauli matrices:
The matrices are needed in order to construct Lorentz invariant quantities and naturally include the spin-orbit coupling in the theory. is the 4-derivative and are the electron spinor fields having as adjoint. In this Lagrangian, only the electrons are treated explicitly as dynamical variables while the nuclei are considered as fixed classical objects following the Born-Oppenheimer approximation.
The dynamics of the free electromagnetic field is described instead by the Lagrangian density ():
where the field tensor :
depends on the 4-vector potential and mediates all the electromagnetic interactions. In the context of polaritonic chemistry, such a term accounts both for the field induced by the electron and the one inherent to the confinement of the molecular system in the cavity. Maxwell’s equations allow to define auxiliary scalar and vector potentials, respectively and . These potentials are not uniquely define, and many potentials lead to the same electric and magnetic field. This is referred to as gauge freedom [44, 46]:
where is a scalar function. is related to the electrostatic component of the electric field, and, in Coulomb gauge, to the longitudinal part of the electric field. On the other hand, is related to the magnetic field, and, in Coulomb gauge, to the transverse component of the electric field. We remind the reader that the electric field can be expressed in terms of scalar and vector potential and the magnetic field in terms of vector potential . The choice of the Lagrangian density for the electromagnetic field is not unique, and different equivalent forms can be used depending on the gauge. Lagrangian 4 is usually the most convenient choice in Coulomb gauge ( [46, 47]). This Lagrangian is not always convenient if other gauges [35, 36] (i.e. Lorenz gauge[46]) need to be used. The light and matter terms are coupled via interaction contributions:
where and are the vector potential and current associated to the field generated by the nuclei. The field in this last contribution to Eq. 8 due to the Born-Oppenheimer approximation is treated as static quantities. The 4-current in Eq. 8 can be expressed as a function of and as:
Finally, the complete QED Lagrangian takes the form:
Eq. 10 will be the starting point for the development of the Hamiltonian formalism derived in the following.
II.2 Hamiltonian formulation in Coulomb gauge
Starting from Lagrangian 10 a Hamiltonian formulation of the theory can be derived performing a Legendre transform:
where the conjugate momenta are given by:
Substituting Eq. 10 in Eq. 11 and using the Green theorem:
where the surface integral is zero, we obtain the following Hamiltonian:
In Eq. 10 no gauge choice has been applied so far. In quantum chemistry application, the Coulomb gauge () is usually a natural choice. This choice allows splitting the electric field into a longitudinal and a transversal component.
In Coulomb Gauge, the longitudinal part of the field only depend on . Under these condition the Gauss law for the electric field ( where is the electron density and is the vacuum permittivity) becomes the Poisson’s equation () and therefore:
Collecting this term with in Eq II.2, gives the well known electron-electron Coulomb repulsive contribution .
To facilitate the quantization (see section II.2.2), it is usually convenient expressing Hamiltonian II.2 in terms of the so-called normal variables. These variable are defined in a way that is quite similar to the ”ladder operator” method used to solve the quantum harmonic oscillator [48, 44]. To begin with, we rewrite Maxwell’s equation in reciprocal space in the following way:
From the previous equation one notices the following relationship when :
where and . From the previous equation, it appears natural to introduce two new variables, even if :
Using the previous relationship one is able to express the electric and magnetic field in terms of normal variables:
and can be obtained using a Fourier transform. Using these normal variables, the Parseval-Plancherel identity [44] allows writing the following:
and changing in the second term of the right-hand side of Eq.21 we can rewrite the electromagnetic field Hamiltonian as:
where the short notation has been used.
The vector potential can be expressed in terms of the normal variables
Substituting Eq.s 25 and 26 in Eq. II.2 we can obtain the following expression for the classical Hamiltonian:
II.2.1 Retardation effects and other QED corrections
In relativistic quantum chemistry, a lot of efforts have been devoted to find effective strategies to include radiative QED corrections [49, 50, 51] (retardation effects, vacuum polarization, Lamb-Shift, etc.), in particular with Bound State QED techniques (BSQED) [52]. Even though such contributions to the electronic properties of the system are fairly small and are in general neglected when performing standard relativistic quantum chemistry calculations, they become quite sizeable when compared with polaritonic effects and should therefore be included.
QED corrections are due to the electromagnetic field-matter interaction term , where . This interaction term contains all the Feynmann graphs related to the various QED corrections.
Such graphs involve what are called virtual-photons, they are internal photonic lines in these graphs and do not correspond to a real particle that can be detected. However, in the formalism of QFT, these virtual photons are the one mediating the electromagnetic interaction between particles. On the other hand, the same interaction term account for the contribution of processes involving real photons. In this case, they appear as external lines in the Feynmann graphs, and they correspond to the particles which are detected in experiments. In BSQED, one is in general interested in the graphs involving virtual photons, whereas in polaritonic chemistry the focus is on the graphs involving real photons. In relativistic polaritonic chemistry, since we cannot a priori neglect radiative QED effects, we need a strategy to include such effects. One strategy could be to adopt a formalism similar to the one of BSQED, however, such formalism, while providing remarkable accuracy for atomics systems, is not really applicable to molecular electronic structure calculations. Fortunately, it is possible to include lower order QED radiative corrections using effective potentials (Breit contribution, Uehling [32, 53, 45, 54], Wichmann-Kroll [55, 54] potential and Self-Energy contribution [37, 38, 45, 54]). Hence, in this way, the dominant graphs involving virtual photons are accounted for, and the formalism of QED is used to describe processes involving real-photons. This strategy remains reasonable since it exploits the linear nature of the Dirac-Hamiltonian and of the interaction term, preserving the additivity of the energy contributions. We would like to remind the reader that some of these contributions include virtual electron-positron pairs. Therefore, using strategies based on effective potentials commonly used in relativistic quantum chemistry are not applicable beyond the no-pair approximation. [54, 56].
Among the radiative QED corrections, the largest term is the so-called Breit term, which is related to retardation effects [53, 45, 57]. Physically, this term originates from the time needed to mediate the electromagnetic interaction between particles, and as the velocity of such particles get closer to the speed of light, such effect becomes more significant. Even though such term contributes to minor energy differences, it can be quite significant for phenomena involving excitation from the core-orbitals, and the effect of such a contribution to the electronic molecular orbital energies has been investigated by Aucar and coworkers [58]. In this work, the possible competition between radiative QED corrections (in particular the Breit contribution) and polaritonic effect is investigated (see Result section III).
There are many ways to derive the Breit term, but one simple way is to work within the Lorenz gauge () which has the advantage of yielding explicitly retarded scalar and vector potentials as done in Ref.s [45]. resulting in an expression of the Full-Breit term as:
which is the retarded electromagnetic interaction without any approximation. For an explicit derivation of this term we refer to Ref.s [53, 45, 54] Since this energy contribution is frequency dependent, it is in general expanded around to obtain the full frequency independent Breit term. Therefore, in relativistic quantum chemistry one generally uses the following 2-body Full-Breit potential:
The first term contains a magnetic contribution due to the transverse component of the vector potential known as the Gaunt term. The second one is related to the retardation effects and is called the Breit term. The 2-body Breit potential includes both these contributions.
In this work, additionally to the Coulomb 2-body potential, we considered the following 2-body operator known as the Coulomb-Breit operator:
which includes both Coulomb and Breit 2-body potentials. In the scientific community, it is fairly accepted to include the Full-Breit term at the HF level in the no-pair approximation. However, using this approach with more refined models (e.g. correlated methods, etc.) is still strongly debated in the field [54].
II.2.2 Hamiltonian Quantization
A quantized form of Hamiltonian 27 can be obtained by promoting the normal variables (/) to the corresponding -dependent field operators (/):
satisfying the following commutation relations:
The explicit time-dependence present in the normal variables has been removed in the field operators expressed in the Schrödinger picture. In terms of the field operator the vector potential becomes:
In the confined space of the optical cavity the wave vector assumes discrete values consequently defining a discrete spectrum of field modes characterized by the direction of the vector and by the polarization of the field oscillations (). In the present case, the discrete vector potential only accounts for real photons, the other contributions due to the transverse part of vector potential (Gaunt, Breit, etc.) could be included as effective potentials (or more elaborate methods based on QED Hamiltonians [52, 49, 50]). In particular, in this work, the Gaunt and Breit contribution are included in the 2-body potential in Eq. 30. Since the molecular system we aim to investigate are usually significantly smaller than the wavelength of the cavity field, we are entitled to apply the dipole approximation imposing that and consequently that in Eq. II.2.2 is equal to 1. This will significantly simplify the photonic part of the Hamiltonian.
Also for the fermionic part, the spinor fields are promoted to spinor fields operators satisfying the following equal-time anti-commutation relations:
The QED Hamiltonian can then be written as:
Where
Here the index collects both the mode number and its polarization.
II.2.3 Relativistic Pauli-Fierz Hamiltonian in the length gauge
When we deal with molecular systems it is usually more convenient to apply a unitary transformation to the field modes, allowing for a direct coupling between the field operators and the molecular dipole. This transformation take the name of length-gauge transformation:
where . Hamiltonian II.2.2 can be transformed in the length-gauge form by application of Eq. 37:
followed by a rotation of the photonic coordinates . Transformation in Eq.38 will induce a cancellation of the term due to the change in the momentum. The light-matter coupling is now related to the molecular dipole. For details regarding the applications of these transformations, refer to Appendix A and to Ref. [59]. The final Hamiltonian in length gauge appears as:
This Hamiltonian has now a shape that is equivalent to the standard Pauli-Fierz Hamiltonian usually applied in non-relativistic polaritonic chemistry, where molecular orbitals are replaced by molecular spinors. Equation II.2.3 has an apparent origin dependence, coming from the presence of the dipole operator . This problem can be solved by the coherent state transformation with [24]. Applying this transformation to the RPF Hamiltonian, one then can use the coherent state basis for the photonic states and is left with the following Hamiltonian:
In Eq. II.2.3 the presence of the expectation value of the total dipole moments ensures that the Hamiltonian will not explicitly depend on the origin of the reference system. In atomic and molecular physics/chemistry, it is usually more convenient to expand the electronic field on a basis of atomic/molecular orbitals which are solutions of the Dirac equation in an external potential (the Coulomb potential of the nuclei). The field operator expanded on such a basis writes as:
where are the atomic solutions. Notice that are atomic spinors, and therefore correspond to both its conjugate and transpose. In order to avoid confusion, the transpose symbol has been omitted. The electronic creation () and annihilation () satisfying anti-commutation rules:
Using these operators the energy components of Hamiltonian II.2.3 can be expressed in a second quantized form implementable in a quantum chemistry code. In the following, we will use a shorthand notation for the mono and bi electronic integrals that is defined as follows:
where is a generic one-body operator and are the well known two-electrons integrals. Therefore, the second quantized (sq) Hamiltonian is:
where:
It is important to notice that Hamiltonian 46 allows both for positive and negative energy states. This point will be discussed in details in the next section (II.2.4).
II.2.4 Treatment of the negative energy states
Physically, the negative energy solutions correspond to positronic states. In relativistic quantum chemistry, the negative energy states are usually filled with electrons, therefore, no electron can fall in these states. However, if one electron is removed from the negative energy states, it would leave a hole corresponding to the creation of a positively charged particle, the positron. This reinterpretation becomes clearer if we split the sum over all state into a sum over the positive energy states and another over the negative energy ones.In this picture, the negative energy electron creation (annihilation) operators are reinterpreted as positive energy positron annihilation (creation) operators (see Ref.s [60, 61, 62, 56]). In this framework, the field operator can be decomposed in this way:
where are respectively positron annihilation and creation operators. Using these operators, Hamiltonian 46 can be rewritten in terms of both electronic and positronic contributions:
where:
and and the modified one and two-electrons integrals have been defined in Eq.47. Barred indices correspond to positronic indices.
II.3 The Hartree-Fock approximation
Since it treats explicitly all the interactions between relativistic electrons/positrons and the photons of the cavity field, Hamiltonian 51 represents the perfect starting point for the development of ab initio theories that can be used to simulated molecular systems containing heavy atoms in optical cavities. In this section, we will use Hamiltonian 51 to develop the first relativistic 4-components Hartree-Fock (HF) approach for polaritonic chemistry. In quantum chemistry, HF represents the simplest approximation respecting the right symmetry of all the involved particles. HF also provides access to physically meaningful sets of atomic/molecular orbitals that can be subsequently used to develop more accurate correlated theories. HF is based on the independent particle approximation, meaning that the involved quantum species do not directly interact between each other but travels in space under the action of the average potential of the other particles (mean field approximation). For the purely electronic case, this implies that the wave function is described by a single Slater determinant of atomic/molecular orbitals or spinors in the relativistic case:
This ansatz ensures that the anti-symmetry of the wave function for a particle exchange will be respected. In this case we deal with three kinds of particles: electrons, positrons and photons, each one respecting its own statistics (fermionic for electrons and positrons and bosonic for photons). The independent particles condition, in this situation, can be enforced defining the wave function ansatz as the product of distinct wave functions for every species:
where and are single determinants for electronic and positronic spinors respectively, is the vacuum state associated to the photonic mode . A similar ansatz has been already applied in the non-relativistic version of the QED-HF approach, and its implications are discussed in detail in Ref. [24]. Ansatz 55 can be rewritten, for practical reasons, in terms of second quantized electronic and positronic operators as:
where and are the electronic and positronic vacuum states.
where :
Notice that at the HF level, all terms that do not conserve the number of particles will have a zero expectation value. This is the case for the so-called bilinear term in Eq.s 46 and 51 which involves only one annihilation (creation) operator and therefore changes the number of photon. Consequently, this term gives a zero contribution to the ground state energy.
II.3.1 The Fock-operator
As for the bare electrons case, the HF problem can be solved by minimizing Eq. II.3 for variations in the spinors. This can be done by performing a rotation of the atomic orbitals using the following operator:
and then imposing the stationary condition:
From this minimization we can obtain, following standard orbital rotation techniques [63, 57, 61], a new Fock operator with matrix elements:
where are either positron or electron indices. Hence the Fock operator in block-matrix form is the following:
where indices () correspond to occupied(unoccupied) electronic spinors. Barred indices () refer instead to positronics orbitals. The Fock operator in Eq. 68 can be applied in a Roothan-Hall like procedure [63]:
to optimize the orbital coefficients. It is important to highlight that exactly as in the non-relativistic QED-HF approach (see Ref. [24] the Fock matrix in Eq. 68 suffers from an explicit origin dependence if charged systems need to be investigated. This problem can be solved applying the so-called Strong Coupling (SC) HF approximation proposed in Ref. [27]. In this paper, we will focus, for the moment, on neutral molecular system and an SC extension of our method will be the topic of a future work.
In standard relativistic quantum chemistry, the positrons are usually removed from the treatment by application of the no-pair approximation [57, 45]. Consequently, the spinor optimization procedure is in fact an excited state problem where the ground state energy corresponds to the first positive energy state. This is referred to as a minimax problem [57, 45]. In our approach, equivalent results can be obtained when no explicit positron is included in the wave function ansatz:
consequently, the projection of Hamiltonian 51 on such wave functions will cancel all positronic dependent terms in Eq. II.3.
This ansatz is the natural extension of the one usually used in non-relativistic QED-HF [24, 27]. In this paper, this simplified ansatz will be used to generate the results presented in section III. The positronic degrees of freedom will be included instead explicitly in a future implementation of the method.
II.3.2 Kinetic Balance in presence of the field
In relativistic quantum chemistry, without including positrons, the spinors can be expressed as two-components objects:
where and are called respectively ”Large” and ”Small” components. In practice, the spinor solutions are expanded on a basis-set: where and is a 2-spinor. In order to avoid the so-called varitional collapse of the solution, a constraint on the large and small components of the basis set is applied [53, 45]. In the presence of a vector potential, such constraint has the following form:
where is the vector of the three Pauli matrices and is the generalized momentum. Such prescription is called the magnetic balance condition [64, 65, 66, 67]. In standard relativistic quantum chemistry, the momentum is simply . We refer to this condition as the kinetic balance [53, 45]. In our case, since the generalized momentum is a priori depending on the field, the magnetic balance needs to be in principle satisfied. This could require to have a field-dependent small-component. However, in the present context, we have shown that in the dipole-approximation, applying the length gauge transformation, the momentum of the theory is therefore transformed in the following way (see App.A):
Consequently, the kinetic balance condition:
can be applied in this case. It is important to stress that this is only possible under the dipole-approximation, otherwise, the length gauge transformation would not provide such a simple shape for the Hamiltonian and the associated momentum.
II.4 Excited states properties
In polaritonic chemistry, the signature property emerging from the strong coupling condition is the Rabi-splitting. It represents the energy separations between the polaritons formed by the mixing between matter and field states (see Fig. 2).
Calculating Rabi splittings, which could directly be compared with experimental data, requires access to the excited states of the coupled light matter system. At the HF level they can be simulated recurring to linear response theory. Recently, Castagnola et al. [68] presented an HF linear response theory for non-relativistic polaritonic systems. In this section, we present an extension of this approach to 4-components Dirac HF.Using linear response theory, excitations energies () can be obtained by solving the well known Casida equation [69]:
The polaritonic espressions of the and matrices as derived in Ref. [68] assumes the form:
where and have matrix elements:
A similar approach, has been recently proposed by Konecny et al. [30] to calculate excitation energies at the QED Dirac Kohn Sham level of theory.
III Results
All the results presented in this section have been obtained from the development version of the PySCF software package [70, 71]. In particular, the Pol-DHF approach has been implemented by applying modifications to the integrals used in standard DHF [72]. So far, to simplify the interpretation of the results, we neglected the positrons in the treatment. A detailed discussion of the effects induced by the explicit inclusion of the positronic degrees of freedom will the topic of a future follow-up paper. All calculations have been performed using the all-electron x2c-SVP basis-set [73, 74, 75, 76] for the large component. The small component has been obtained via the kinetic balance prescription. The coupling value has been set to 0.05 a.u..
III.1 Ground state properties
In this section, we analyze the field induced effects on the ground state properties of three metal hydrides (CuH, AgH and AuH). These complexes contain metals belonging to different periods of the 11th group of the periodic table. Going down the groups, the velocity of the electrons (in particular of the inner ones) increases, approaching finite fractions of the speed of light in the gold case. In fact, gold complexes are well known to have peculiar electronic properties due to the significant relativistic effects [77, 78, 79, 80]. In this paper, we will analyze how the polaritonic effects compete with the relativistic ones generating modification of the molecular electronic structure. In Table 1 we compare the total energies calculated at the DHF level of theory with those evaluated including the field (Pol-DHF).
molecule | DHF (Hartree) | Pol-DHF (Hartree) | E (eV) |
CuH | -1653.11275 | -1653.11843 | -0.1543 |
AgH | -5338.68917 | -5338.69790 | -0.2374 |
AuH | -21639.06979 | -21639.07709 | -0.1988 |
The field induced effects, despite small compared to the total energies, represent still a sizable (some tens of eV) variation on the energy of the system. In particular, it is interesting noticing that the field induced energy variation is quite similar for copper and gold (Cu - 0.15 eV, Au - 0.20 eV) while it is slightly bigger (in absolute value) for silver (Ag - 0.24 eV). This trend could be actually explained if we realize that under the QED-HF approximation, the only field term contributing to the total energy of the system is the dipole self-energy contribution . As highlighted in Eq. 47 the one-electron term coming from the dipole self-energy directly depends on the molecular quadrupole, and it is well known from the literature (see Ref. [81]) that this as many other electronic properties (atomic radius, etc.) [82] show a very similar trend if we move down the group.
The effects observed on the total energy values are an indirect observation of the variations induced by the field on the system’s molecular orbitals (MO).
In Figure 3 we compare the orbital energies for the valence orbitals ( and ) of the three complexes evaluated at the Pol-DHF level with those calculated at the DHF and scalar 2-components X2C level of theory [83, 70, 71]. As expected, the biggest variations in the orbital energies can be observed comparing the scalar X2C results with the DHF ones. in this case, the usual destabilization of the orbitals and consequent stabilization of the MOs can be observed. The reduction of the gap between these occupied MOs decreases with the increasing relativistic character of the metal (smaller for Au than Cu). Much smaller, even if sizable, is the effect produced by the field.
molecule | (meV) | (%) |
CuH | 26.9 | 0.62 |
AgH | 38.4 | 0.74 |
AuH | 11.2 | 0.98 |
The field induce effects () on the energy gap, barely appreciable in Figure 3, are reported in Table 2. As already observed for the total energy also in the field induced variation of the energy gap is larger for AgH compared to the CuH and, in particular, to AuH. It is crucial, to point out that also in the gold case (smallest effect) the cavity field is able to induce a 10 meV variation on the orbital gap. These values, represent a reasonable fraction ( or higher) of a kcal/mol, relevant to observe variations of the chemical properties. Even more interesting is observing that the relative change induced by the field () is larger for AuH (about 1% of the total gap) compared to the other systems (0.7% for AgH and 0.6% for CuH). Notice that these field induced variations of the energy gap bring (for all the systems) to a reduction of the gap due to a very small stabilization of the MOs accompanied by a larger destabilization of the orbitals.
III.1.1 Gaunt, Breit and polaritonic contributions to ground state energies
As mentioned in section II.2.1 the radiative QED corrections and in particular the retardation effects can compete or sum up to the polaritonic contributions in a relativistic molecular system. Here we perform a detailed comparison between the energy variations introduced by quantum field, with the Gaunt and Breit terms contributing to the full retarded potential.
molecule | No Pol | Pol | ||
Gaunt (eV) | Breit (eV) | Gaunt (eV) | Breit (eV) | |
CuH | 20.6381 | -1.8419 | 20.6381 | -1.8419 |
AgH | 109.7382 | -10.8607 | 109.7380 | -10.8607 |
AuH | 544.0871 | -66.5148 | 544.0874 | -66.5149 |
In Table 3 the Gaunt/Breit corrections to the DHF and Pol-DHF energies are presented. These effects are quite sizable, in particular for AuH that is the system showing the largest relativistic effects. In this case the trend is monotonic and both the Gaunt and Breit effect increase going down the group of the periodic table. It is important to notice that if the absolute energy variation is taken into account, these effects are at least one order of magnitude larger than the effects generated by the quantum field. Emblematic is the comparison between the Breit correction in CuH (the less relativistic system) and the corresponding polaritonic energy correction, shown in Table 1. These observations clearly indicate that, for a molecular system coupled to a quantum field, attention needs to be paid in omitting the retardation effects, a generic pratice in standard electronic relativistic quantum chemistry simultations.More striking results can be obtained if we perform the same analysis on the orbital energies. In Figure 4 the energy contributions (in logaritmic scale) due to the Gaunt, Breit and polaritonic terms on the energies of the occupied MOs are presented for the three systems.
The most evident aspect is the very wide variation range of the energy contributions for different molecular orbitals. In general, the effects due to the Full-Breit term are much more sizable for the core orbitals, while they monotonically decrease moving toward the valence ones. This trend is clearly expected, since the inner orbitals host the fastest electrons that are more affected by the relativistic effects. This behavior is obviously more evident for AuH than for CuH. The polaritonic contribution shows instead much smaller variations, and the effect slightly increases moving toward the valence. In the core region, the retardation effects are always larger, while in the valence region the polaritonic effects are dominant. It is interesting highlighting that for the intermediate orbitals all the effects are quite comparable, confirming that the inclusion of retardation effects could produce crucial effects in molecular systems coupled to photons. There are anyway cases where, neglecting such effects, as usually done in relativistic quantum chemistry, is still reasonable. For instance, if we are interested in excited states properties (i.e. Rabi Splitting, etc.) involving only valence electrons, a phenomenon recently analyzed by Konecny et al. in Ref. [30], then omitting retardation effects in the treatment is clearly acceptable. However, if excitations from inner orbitals (i.e. core excitations, etc.) or magnetic properties of the system need to be analyzed the inclusion of the retardation effects will become a crucial point in the treatment.
It is important to point out that the implications of the effects presented in this section are, in some sense, minimized by the absence of electron-electron and electron-photon correlation in the treatment. For instance, the HF approximation removes the frequency dependence of the ground state energy from the frequency of the field. This dependence can be recovered only by including correlation into the model. We expect, that the inclusion of frequency dependent terms, will be crucial in particular to describe resonant processes. The inclusion of electron-electron and electron-photon correlation will be the main topic of a future follow-up paper.
III.2 Excited state properties
In the section III.1, we focused on the field induced effects on the ground state properties of three metal hydrides of the 11th group of the periodic table. In this section, we analyze in details the effects generated by the photons on the optical properties of the AuH complex [84, 85]. Similar analysis for CuH and AgH is presented in Appendix B. In this system, because of strong relativistic effects, we expect to observe spectra very different compared to those that can be simulated without taking relativity into account. This is evident if we look at Figure 5.
As it can be seen, at the HF level, the triplet state () is found to be higher in energy than the singlets states, in disagreement with the experimental data from Ref.s [84, 85]. The inclusion of scalar relativistic correction partially resolve this issue, but it underestimates the energy by about 0.45 eV compared to DHF calculations. DHF does not only reproduce the right ordering of the states, but it also describes the breaking of the degeneracy to form (from the triplets) the and states. For an improved readability of the plot Figure 5 the Pol-DHF data have been calculated off-resonance. This choice highlights the energy shift due to the dipole self energy. In resonance conditions also the Rabi splitting (discussed later in Figure 6) would be visible.
Comparing this data with the bare electronic DHF ones, we see that the field (in off resonance conditions) induces a stabilization of the and states but without inducing any change in the ordering of the states.
Finally, in Fig.6 we have reported the dispersion of the AuH excitation energies as a function of the cavity frequency. The excitations falling in the investigated energy range, refers to the and states obtained by the splitting of the triplets. For a non-relativistic system we should not observe any Rabi splitting due to the selection rule. In this case instead, due to the strong spin-orbit coupling, a sizable splitting can be observed at the crossing between the first photonic replica of the ground state () with the state. For this system, the energy difference between and , associated to spin-orbit coupling, is significantly larger (0.04 eV) than the Rabi-splitting.
This spin-orbit coupling induced singlet-triplet rabi-splitting was already reported for a different system by Koneckny et al. in Ref. [30]. These observations clearly prove how the electromagnetic field can be used to manipulate and control inter-system crossing processes and consequently the phosphorescence of complexes containing heavy atoms. In App. B a similar discussion is reported also for CuH and AgH.
IV Conclusions
In this paper we proposed a reformulation of relativistic QED allowing for an easier development of ab initio methodologies to simulate heavy atoms molecular complexes in strong coupling conditions. Using this theoretical ground, we reported the development and implementation of the first relativistic polaritonic wave function based ab initio method, namely Pol-DHF. The theory has been presented starting from the standard Lagrangian and has been derived into a usable implementation of the Pol-DHF code. Considering the possible competition with the polaritonic effects, the radiative QED corrections have been explicitly included into the treatment. After providing our roadmap to the implementation, we presented applications of Pol-DHF to three metal hydrides: CuH, AgH and, AuH. These systems were an excellent test case in order to assess the magnitude of the polaritonic effects in comparison to the relativistic effects. To do so, we evaluated the influence of the polaritonic effects on the gap for the three systems. The polaritonic effects resulted having the largest relative influence on AuH for which relativistic effects are more prominent. Afterwards, we provided a detailed analysis of the competition between the Gaunt, Breit and polaritonic effects on the ground state and orbital energies. We have shown that even though the Full-Breit term represents a significantly larger contribution to the total ground state energy, its effect is extremely orbital dependent. In particular, it is much larger on the core orbitals compare to the valence, where instead the polaritonic effects dominate. Therefore, we could conclude that neglecting the Full-Breit term (practice usually applied in relativistic quantum chemistry) in strong coupling conditions is only possible if properties involving valence orbitals need to be investigated. Lastly, we presented excited state calculations for AuH at the TD-Pol-DHF level. As already observed by Koneckny et al. [30], we have shown that using a fully relativistic polaritonic theory the appearance of Rabi-Splittings at the crossing between singlet and triplet potential energy surfaces can be observed. These results clearly open new possibilities for the control of the inter-system crossing in photochemistry. We strongly believe that the methodologies and the applications presented in this paper will represent a groundbreaking step toward the simulation of relativistic molecular systems strongly coupled to photons, field that is recently finding many interesting applications in photochemistry but also in spintronics and quantum computation.
Acknowledgements.
Guillaume Thiam, Riccardo Rossi and Enrico Ronca acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Program (Grant ERC-StG-2021-101040197-QED-SPIN). Henrik Koch acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (Grant Agreement 101020016).Appendix A Length gauge transformation
In section II.2.3, we present the results of the application of the length gauge transformation. In this appendix, we provide a detailed derivation of such results. Since most of the terms of Hamiltonian II.2.2 commute with the operator, only two terms get modified, the one involving the momentum and the one involving the photon number operator . Using the special case of the Becker-Campbell-Hausdorff formula:
Let us consider the commutator only:
and therefore
It is important to emphasize that the dipole approximation strongly simplifies the expression of the terms involving the momentum. The other noticeable fact is that the terms involving cancels out (the translated momentum bring out a term cancelling out with the unmodified term).
Let us now look at the modification of the photon number operator. The Becker-Campbell-Hausdorff formula implies that:
The first commutator gives:
where . The second commutator gives:
Therefore, the Hamiltonian then reads:
Appendix B Excited states properties of CuH and AgH
In Fig. 7, we report the dispersion of the CuH excitation energies with respect to the cavity frequency. Also in this case, despite significantly smaller than the one observed for AuH in Fig. 6, we detect an observable Rabi-splitting. This is expected since the Cu atom is much lighter than gold and relativistic effects, including the spin-orbit coupling, are significantly smaller. In this case, the energy difference between the and state is 20 times smaller than for AuH ( 0.002 eV).
Similar results can be observed also in Fig. 8 for AgH. In this case, the observed behavior is somehow intermediate between the one of CuH and the one of AuH, consistently with the fact that Ag is heavier than Cu but lighter than Au. For this system the degeneracy between the three states is slightly lifted due to the 4-component treatment. However, contrary to AuH the energy difference between the and state is significantly smaller ( 0.009 eV). Notice that for both CuH and AgH, the Rabi-splitting is larger than such energy difference.

### v2

@@
-A comprehensive theory for relativistic polaritonic chemistry: a way toward a four component ab initio treatment of molecular systems
+A comprehensive theory for relativistic polaritonic chemistry: a four components ab initio treatment of molecular systems coupled to quantum fields
 Abstract
@@
 The relativistic quantum electrodynamics (QED) theory has been rewritten with a formalism close to relativistic quantum chemistry.
-This general framework represents the ideal starting point to extend the main quantum chemistry methods to relativistic QED.
+This general framework represents the ideal starting point to extend the main quantum chemistry methods to relativistic polaritonic.
 The Polaritonic Dirac Hartree Fock (Pol-DHF) approach is the first method we propose based on this theory.
@@
 The method is able to include not only the effects of the photons but can be easily extended also to include explicit interactions with positrons.
-Application of Pol-DHF to three metal hydrides revealed the importance of including radiative QED corrections to the treatment in strong coupling conditions.
+Application of Pol-DHF to three metal hydrides revealed the importance of including relativistic effects to the treatment in strong coupling conditions.
 Due to an accurate description of spin-orbit coupling, the method is able to reproduce polaritonic effects happening at the crossing between singlet and triplet potential energy surfaces.
@@
 The strong coupling condition is usually reached inside properly designed optical devices.
-The simplest example is the Fabry-Pérot cavity [6] (Fig 1), made of two highly reflective planar mirrors, that confine the photons in a small quantization volume and leading to a significant enhancement of the light-matter coupling.
-The field properties and therefore the polaritons properties can be controlled changing the geometry and the materials of the cavity.
-Several demonstrations of polaritonic effects on different physical properties such as absorption spectra, photochemical reaction rates, and conductivity have already been observed in the experiments [2, 7, 3, 5, 8, 9, 10].
-Probably the most famous demonstration was obtained, in the experiments performed by the Ebbesen’s group in Strasbourg.
-In particular, they demonstrated that strong coupling to molecular vibrations can be used to catalyze, slow down or even induce selectivity in chemical reactions [11, 5, 12].
+The simplest example is the Fabry-Pérot cavity [6] (Fig 1), made of two highly reflective planar mirrors, that confine the photons, leading to a significant enhancement of the light-matter coupling.
+The cavity frequency, and therefore, the polaritons properties can be controlled changing the geometry and the materials in the device.
+Several manifestations of polaritonic effects on different physical properties such as absorption spectra, photochemical reaction rates, and conductivity have already been observed in the experiments [2, 7, 3, 5, 8, 9, 10].
+What is probably the most well-known demonstration of such effects was obtained in the experiments performed by Ebbesen’s group in Strasbourg.
+In particular, they evidenced that strong coupling to molecular vibrations can be used to catalyze, slow down or even induce selectivity in chemical reactions [11, 5, 12].
 These observations opened a new field that is now known as polaritonic chemistry [13].
 In these experiments, the photonic states are usually coupled either to electronic or vibrational states of the molecular systems.
-However, the electromagnetic nature of the field allows also for modifications of their magnetic properties if a coupling to the spin states is exploited.
+However, the electromagnetic nature of the field also allows for modifications of their magnetic properties if one exploits the coupling to spin-states.
 In this way, a fine control of the magnetizability and aromatic properties of molecules [14], of spin qubits [15, 16] and of spin phases of materials [17, 18] can be obtained.
@@
 The problem can be circumvented using planar superconducting devices commonly used in Circuit-QED experiments [19].
-Using similar devices Affronte’s group has been able to manipulate the spin properties of a Single-Molecule Magnet [16, 20].
-Such an accomplishment opened large possibilities of application in spin qubits based quantum computation.
+Using similar devices, Affronte’s group has been able to manipulate the spin properties of a Single-Molecule Magnet [16, 20].
+Such an accomplishment unlocked many potential applications in spin qubits based quantum computation.
 Despite the many improvements in the fabrication of more effective optical devices and the impressive accuracy reached [21, 22] by polaritonic chemistry experiments, many fundamental aspects still remain to be understood.
@@
 In recent years many ab initio methods, able to treat electron-electron and electron-photon correlation at different level of accuracy have been developed [23, 24, 25, 26, 27, 28].
-However, they have been mainly applied to investigate cavity induced effects on the electronic and vibrational degrees of freedom of molecular systems.
+However, they have mainly been applied to investigate cavity induced effects on the electronic and vibrational degrees of freedom of molecular systems.
 Only very recently, Barlini et al.
@@
 Spin and consequently magnetic properties are intrinsic relativistic features of molecules.
-An accurate investigation of these properties requires then inclusion of relativistic effects at different level of accuracy.
-This becomes particularly crucial when, like in Affronte’s experiments [16, 20] Single-Molecule Magnets containing lanthanides atoms need to be used to have sufficiently long living magnetizations.
-Moreover, as already discussed in Refs [28, 29] for Van-der Waals interactions, the field has sometimes to power to enhance small effects that are usually neglected in absence of the photons.
+An accurate investigation of these properties then requires the inclusion of relativistic effects at different level of accuracy.
+This becomes particularly crucial when, like in Affronte’s experiments [16, 20, 29, 30, 31, 32, 33] Single-Molecule Magnets containing lanthanides atoms need to be used to have sufficiently long lasting magnetizations.
+Moreover, as already discussed in Refs [28, 34] for Van-der Waals interactions, the field can sometimes enhance small effects that are usually negligible in the absence of photons.
 From these considerations rises the emergency of a consistent relativistic quantum electrodynamical ab initio theory able to include all the necessary effects.
@@
 [35] proposed a relativistic response theory to investigate electronic excitations of relativistic molecules in optical cavities.
-Despite this last approach was able to capture interesting effects like cavity-induced singlet-triplet interactions, it is neglecting additional photon driven effects that might be crucial if magnetic properties want to be investigated.
-Moreover, this approach does not account for modifications induced by the field on the system’s ground state.
+Despite this last approach being able to capture interesting effects like cavity-induced singlet-triplet interactions, it is neglecting additional photon driven effects that might be crucial if magnetic properties want to be investigated.
+Moreover, at the moment, the implemented approach does not account for modifications induced by the field on the system’s ground state.
 This is a crucial aspect if we are interested to cavity driven effects on the core properties of relativistic atoms.
-In this paper, we develop the first fully relativistic ab initio method to describe the ground state of molecular systems coupled to photons in optical cavities.
-We start the development from the most general Quantum Electrodynamical Lagrangian formulation to then propose a new and totally method independent approach to develop relativistic ab initio theories for integrating electron-photon systems.
+In this paper, we develop the first fully relativistic wave-function based ab initio method to describe the ground state of molecular systems coupled to photons in optical cavities.
+We start the development from the most general Quantum Electrodynamics Lagrangian to then propose a new and totally method independent approach to develop relativistic ab initio theories for integrating electron-photon systems.
 This formalism represents the starting point for the development of mean-field but also correlated methodologies.
@@
 renormalization of the electron mass [40, 41]).
-Their inclusion in many-electrons systems is usually carried out using effective QED potentials.[37, 38, 39, 40, 41].
+Their inclusion in many-electrons systems is usually carried out using effective QED potentials [42, 43, 44, 45, 46], however, strong efforts are being deployed to accurately include this corrections using the Bound State QED formalism [47, 48, 49].
 In a recent paper [50], it has been demonstrated that Breit contribution and these QED corrections effects on the Ionization Potential (IP) and Electron Affinities (EA) of gold are of the same order of magnitude (0.01-0.03 eV).
@@
 This is a crucial point that will be discussed in this paper and that motivates the proposed general reformulation of the relativistic polaritonic framework.
-The paper is structured as follows: in section II a generic derivation of a relativistic QED Hamiltonian theory will be presented, starting from the standard QED Lagrangian.
-In this section, the choice of the gauge is discussed in detail and the different energy contributions are analyzed.
+The paper is structured as follows: in section II, a generic derivation of a relativistic QED Hamiltonian theory will be presented, starting from the standard QED Lagrangian.
+In this section, the choice of the gauge is discussed and the different energy contributions are analyzed.
 At this stage, no approximations are used in order to set up a theory that could be the starting point for every quantum chemistry modelization.
-In subsection II.2.2 the coulomb gauge Hamiltonian, after application of the dipole approximation, is used to develop the first Relativistic Polaritonic HF (Pol-DHF) approach.
+In subsection II.2.2, the coulomb gauge Hamiltonian, after application of the dipole approximation, is used to develop the first Relativistic Polaritonic HF (Pol-DHF) approach.
 In this context, large space has been dedicated to strategies used to deal with negative energy states and to improve convergence issues.
-In section III Pol-DHF has been applied to investigate electronic properties of small diatomic molecules containing heavy atoms.
+In section III, Pol-DHF has been applied to investigate electronic properties of small diatomic molecules containing heavy atoms.
 We end the paper with conclusions and perspectives.
@@
 The proposed methodology is then used to develop the first Hartree-Fock (HF) based approach for relativistic molecular systems strongly coupled to quantum fields.
-For convenience reasons, Gaussian units will be used through the whole derivation unless specified otherwise.
+For convenience reasons, Gaussian units will be used during the whole derivation unless specified otherwise.
 II.1 The Quantum Electrodynamics Lagrangian
@@
 The light and matter terms are coupled via interaction contributions:
-where and are the vector potential and current associated to the field generated by the nuclei.
-The field in this last contribution to Eq.
-8 due to the Born-Oppenheimer approximation is treated as static quantities.
-The 4-current in Eq.
-8 can be expressed as a function of and as:
+where is an external potential associated to the nuclei in the Born-Oppenheimer approximation. The 4-current in Eq. 8 can be expressed as a function of and as:
 Finally, the complete QED Lagrangian takes the form:
@@
 To begin with, we rewrite Maxwell’s equation in reciprocal space in the following way:
-From the previous equation one notices the following relationship when :
+From the previous equation, one notices the following relationship when :
 where and . From the previous equation, it appears natural to introduce two new variables, even if :
@@
 The vector potential can be expressed in terms of the normal variables
-Substituting Eq.s 25 and 26 in Eq. II.2 we can obtain the following expression for the classical Hamiltonian:
-II.2.1 Retardation effects and other QED corrections
-In relativistic quantum chemistry, a lot of efforts have been devoted to find effective strategies to include radiative QED corrections [49, 50, 51] (retardation effects, vacuum polarization, Lamb-Shift, etc.), in particular with Bound State QED techniques (BSQED) [52].
+Substituting Eq.s 25 and 26 in Eq. II.2 we obtain the following expression for the classical Hamiltonian:
+II.2.1 Retardation effects and QED corrections
+In relativistic quantum chemistry, a lot of efforts have been devoted to find effective strategies to include additional corrections including retardation effects and radiative QED effects (vacuum polarization, Lamb-Shift, etc.) [48, 57, 49], in particular with Bound State QED techniques (BSQED) [47].
 Even though such contributions to the electronic properties of the system are fairly small and are in general neglected when performing standard relativistic quantum chemistry calculations, they become quite sizeable when compared with polaritonic effects and should therefore be included.
-QED corrections are due to the electromagnetic field-matter interaction term , where . This interaction term contains all the Feynmann graphs related to the various QED corrections.
-Such graphs involve what are called virtual-photons, they are internal photonic lines in these graphs and do not correspond to a real particle that can be detected.
+These additional corrections are due to the electromagnetic field-matter interaction term , where . This interaction term is the one responsible for all these corrections.
+In terms of Feynmann diagrams, the latter are related with graphs involving virtual-photons, internal photonic lines in these graphs that do not correspond to a real particle that can be detected.
 However, in the formalism of QFT, these virtual photons are the one mediating the electromagnetic interaction between particles.
@@
 In this case, they appear as external lines in the Feynmann graphs, and they correspond to the particles which are detected in experiments.
-In BSQED, one is in general interested in the graphs involving virtual photons, whereas in polaritonic chemistry the focus is on the graphs involving real photons.
-In relativistic polaritonic chemistry, since we cannot a priori neglect radiative QED effects, we need a strategy to include such effects.
-One strategy could be to adopt a formalism similar to the one of BSQED, however, such formalism, while providing remarkable accuracy for atomics systems, is not really applicable to molecular electronic structure calculations.
-Fortunately, it is possible to include lower order QED radiative corrections using effective potentials (Breit contribution, Uehling [32, 53, 45, 54], Wichmann-Kroll [55, 54] potential and Self-Energy contribution [37, 38, 45, 54]).
-Hence, in this way, the dominant graphs involving virtual photons are accounted for, and the formalism of QED is used to describe processes involving real-photons.
+In BSQED, one is in general interested in the graphs involving virtual photons, whereas in polaritonic chemistry the focus is on the graphs involving photons as external lines.
+In relativistic polaritonic chemistry, since we cannot a priori neglect these additional corrections, we need a strategy to include them into the treatment.
+One strategy could be adopting a formalism similar to the one used in BSQED.
+However, such formalism, while providing remarkable accuracy for atomics systems, is not really applicable to molecular electronic structure calculations.
+In fact, in polaritonic chemistry people mostly rely on the one-mode approximation.
+In practice, the same approximation will be made in Relativistic Polaritonic chemistry.
+Most of the effects due to the continuum of modes can be approximated including lower order corrections using effective potentials (Breit contribution, Uehling [37, 58, 53, 59], Wichmann-Kroll [60, 59] potential and Self-Energy contributions [42, 43, 53, 59]).
+In this way, the dominant graphs involving virtual photons are included while the QED formalism is used to describe processes involving photons as external lines in the diagrams.
 This strategy remains reasonable since it exploits the linear nature of the Dirac-Hamiltonian and of the interaction term, preserving the additivity of the energy contributions.
-We would like to remind the reader that some of these contributions include virtual electron-positron pairs.
-Therefore, using strategies based on effective potentials commonly used in relativistic quantum chemistry are not applicable beyond the no-pair approximation.
-[54, 56].
-Among the radiative QED corrections, the largest term is the so-called Breit term, which is related to retardation effects [53, 45, 57].
+We would like to remind the reader that some of these contributions include virtual electron-positron pairs and that effective potentials are not applicable beyond the no-pair approximation [59, 61].
+Among all aforementioned corrections, we will focus on the so-called Breit term, which is related to retardation effects [58, 53, 62] and in some cases have a magnitude comparable to the other QED corrections effects [50].
 Physically, this term originates from the time needed to mediate the electromagnetic interaction between particles, and as the velocity of such particles get closer to the speed of light, such effect becomes more significant.
 Even though such term contributes to minor energy differences, it can be quite significant for phenomena involving excitation from the core-orbitals, and the effect of such a contribution to the electronic molecular orbital energies has been investigated by Aucar and coworkers [63].
-In this work, the possible competition between radiative QED corrections (in particular the Breit contribution) and polaritonic effect is investigated (see Result section III).
-There are many ways to derive the Breit term, but one simple way is to work within the Lorenz gauge () which has the advantage of yielding explicitly retarded scalar and vector potentials as done in Ref.s [45].
-resulting in an expression of the Full-Breit term as:
+In this work, the possible competition between the aforementioned additional corrections (in particular the Breit contribution) and polaritonic effect is investigated (see Result section III).
+There are many ways to derive the Breit term, but one simple way is to work within the Lorenz gauge () which has the advantage of yielding explicitly retarded scalar and vector potentials as done in Ref.s [53] resulting in an expression of the full-Breit term as:
 which is the retarded electromagnetic interaction without any approximation.
-For an explicit derivation of this term we refer to Ref.s [53, 45, 54] Since this energy contribution is frequency dependent, it is in general expanded around to obtain the full frequency independent Breit term.
-Therefore, in relativistic quantum chemistry one generally uses the following 2-body Full-Breit potential:
+For an explicit derivation of this term we refer to Ref.s [58, 53, 59].
+This frequency dependent energy contribution is expanded around to obtain the full frequency independent Breit term.
+In relativistic quantum chemistry the following 2-body full-Breit potential:
+is usually applied.
 The first term contains a magnetic contribution due to the transverse component of the vector potential known as the Gaunt term.
@@
 In the confined space of the optical cavity the wave vector assumes discrete values consequently defining a discrete spectrum of field modes characterized by the direction of the vector and by the polarization of the field oscillations ().
-In the present case, the discrete vector potential only accounts for real photons, the other contributions due to the transverse part of vector potential (Gaunt, Breit, etc.) could be included as effective potentials (or more elaborate methods based on QED Hamiltonians [52, 49, 50]).
-In particular, in this work, the Gaunt and Breit contribution are included in the 2-body potential in Eq.
-30.
-Since the molecular system we aim to investigate are usually significantly smaller than the wavelength of the cavity field, we are entitled to apply the dipole approximation imposing that and consequently that in Eq.
+In the present case, the discrete vector potential only accounts for contributions due to the transverse part of vector potential (Gaunt, Breit, etc.) could be included as effective potentials (or more elaborate methods based on QED Hamiltonians [47, 48, 57]).
+In particular, in this work, the Gaunt and Breit contributions are included in the 2-body potential in Eq.
+30 at the DHF level of theory using the no-pair approximation.
+The usual dressed mass of the electron has been used instead of the bare one.
+Since the molecular systems of interest are usually significantly smaller than the wavelength of the cavity field, we are entitled to apply the dipole approximation imposing that and consequently that in Eq.
 II.2.2 is equal to 1.
 This will significantly simplify the photonic part of the Hamiltonian.
-Also for the fermionic part, the spinor fields are promoted to spinor fields operators satisfying the following equal-time anti-commutation relations:
+Furthermore, for the fermionic part, the spinor fields are promoted to spinor fields operators satisfying the following equal-time anti-commutation relations:
 The QED Hamiltonian can then be written as:
@@
 II.2.3 Relativistic Pauli-Fierz Hamiltonian in the length gauge
-When we deal with molecular systems it is usually more convenient to apply a unitary transformation to the field modes, allowing for a direct coupling between the field operators and the molecular dipole.
-This transformation take the name of length-gauge transformation:
+When investigating molecular systems, it is usually more convenient to apply a unitary transformation to the field modes, allowing for a direct coupling between the field operators and the molecular dipole.
+This transformation is known as the length-gauge transformation:
 where . Hamiltonian II.2.2 can be transformed in the length-gauge form by application of Eq. 37:
 followed by a rotation of the photonic coordinates .
-Transformation in Eq.38 will induce a cancellation of the term due to the change in the momentum.
+Transformation in Eq.38 induces a cancellation of the term due to the change in the momentum.
 The light-matter coupling is now related to the molecular dipole.
@@
 The final Hamiltonian in length gauge appears as:
-This Hamiltonian has now a shape that is equivalent to the standard Pauli-Fierz Hamiltonian usually applied in non-relativistic polaritonic chemistry, where molecular orbitals are replaced by molecular spinors.
+where and , have been relabeled to emphasize that, due to the length gauge transformation, the latters are modified, and does not coincide with the number of photons anymore [65].
+This Hamiltonian is the analogous of the standard Pauli-Fierz Hamiltonian usually applied in non-relativistic polaritonic chemistry [23, 66, 24], where molecular orbitals are replaced by molecular spinors.
 Equation II.2.3 has an apparent origin dependence, coming from the presence of the dipole operator .
 This problem can be solved by the coherent state transformation with [24].
-Applying this transformation to the RPF Hamiltonian, one then can use the coherent state basis for the photonic states and is left with the following Hamiltonian:
+Applying this transformation to the Relativistic-Pauli-Fiertz (RPF) Hamiltonian, one then can use the coherent state basis for the photonic states and is left with the following Hamiltonian:
+where is the usual one-electron Dirac Hamiltonian.
 In Eq.
-II.2.3 the presence of the expectation value of the total dipole moments ensures that the Hamiltonian will not explicitly depend on the origin of the reference system.
+II.2.3, the presence of the expectation value of the total dipole moments () ensures that the Hamiltonian will not explicitly depend on the origin of the reference system.
 In atomic and molecular physics/chemistry, it is usually more convenient to expand the electronic field on a basis of atomic/molecular orbitals which are solutions of the Dirac equation in an external potential (the Coulomb potential of the nuclei).
@@
 Using these operators the energy components of Hamiltonian II.2.3 can be expressed in a second quantized form implementable in a quantum chemistry code.
-In the following, we will use a shorthand notation for the mono and bi electronic integrals that is defined as follows:
+In the following, we will use a shorthand notation for the mono- and bi- electronic integrals that are defined as follows:
 where is a generic one-body operator and are the well known two-electrons integrals. Therefore, the second quantized (sq) Hamiltonian is:
 where:
-It is important to notice that Hamiltonian 46 allows both for positive and negative energy states. This point will be discussed in details in the next section (II.2.4).
+where refers to the molecular quadrupole and to the polariton number.
+Notice that Hamiltonian 46 allows both for positive and negative energy states.
+This point will be discussed in details in the next section (II.2.4).
 II.2.4 Treatment of the negative energy states
-Physically, the negative energy solutions correspond to positronic states.
-In relativistic quantum chemistry, the negative energy states are usually filled with electrons, therefore, no electron can fall in these states.
-However, if one electron is removed from the negative energy states, it would leave a hole corresponding to the creation of a positively charged particle, the positron.
-This reinterpretation becomes clearer if we split the sum over all state into a sum over the positive energy states and another over the negative energy ones.In this picture, the negative energy electron creation (annihilation) operators are reinterpreted as positive energy positron annihilation (creation) operators (see Ref.s [60, 61, 62, 56]).
+Physically, the negative energy solutions correspond to positronic states [67].
+This reinterpretation becomes clearer if we split the sum over all state into a sum over the positive energy states and another over the negative energy ones.
+In this picture, the negative energy electron creation (annihilation) operators are reinterpreted as positive energy positron annihilation (creation) operators (see Ref.s [67, 68, 69, 61]).
 In this framework, the field operator can be decomposed in this way:
@@
 where:
-and and the modified one and two-electrons integrals have been defined in Eq.47. Barred indices correspond to positronic indices.
+and and the modified one and two-electrons integrals have been defined in Eq.47.
+Barred indices correspond to positronic indices.
+In this article we do not address the role of the fermionic vacuum that is at the heart of many discussions [68, 69, 61].
+Indeed, the choice of the Hamiltonian and the possible reinterpretation of negative energy states has an influence on the nature of the fermionic vacuum which in turn has some consequence on the expectation value of the Hamiltonian.
+Fortunately, at the Hartree-Fock level there is no real dependence on the vacuum nature (however, it becomes important when developing correlated methods).
 II.3 The Hartree-Fock approximation
-Since it treats explicitly all the interactions between relativistic electrons/positrons and the photons of the cavity field, Hamiltonian 51 represents the perfect starting point for the development of ab initio theories that can be used to simulated molecular systems containing heavy atoms in optical cavities.
+Since it treats explicitly all the interactions between relativistic electrons/positrons and the photons of the cavity field, Hamiltonian 51 represents the perfect starting point for the development of ab initio theories suitable to simulate molecular systems containing heavy atoms in optical cavities.
 In this section, we will use Hamiltonian 51 to develop the first relativistic 4-components Hartree-Fock (HF) approach for polaritonic chemistry.
@@
 For the purely electronic case, this implies that the wave function is described by a single Slater determinant of atomic/molecular orbitals or spinors in the relativistic case:
-This ansatz ensures that the anti-symmetry of the wave function for a particle exchange will be respected.
-In this case we deal with three kinds of particles: electrons, positrons and photons, each one respecting its own statistics (fermionic for electrons and positrons and bosonic for photons).
+This ansatz ensures the anti-symmetry of the wave function under particle exchange.
+In this case, three kinds of particles are involved: electrons, positrons and photons, each one respecting its own statistics (fermionic for electrons and positrons and bosonic for photons).
 The independent particles condition, in this situation, can be enforced defining the wave function ansatz as the product of distinct wave functions for every species:
-where and are single determinants for electronic and positronic spinors respectively, is the vacuum state associated to the photonic mode .
+where and are single determinants for electronic and positronic spinors respectively, is the vacuum state associated to the photonic mode and is a shorthand notation for .
 A similar ansatz has been already applied in the non-relativistic version of the QED-HF approach, and its implications are discussed in detail in Ref.
@@
 Notice that at the HF level, all terms that do not conserve the number of particles will have a zero expectation value.
-This is the case for the so-called bilinear term in Eq.s 46 and 51 which involves only one annihilation (creation) operator and therefore changes the number of photon.
+This is the case for the so-called bilinear term in Eq.s II.2.3, 46 and 51 which involves only one annihilation (creation) operator and therefore changes the number of photon.
 Consequently, this term gives a zero contribution to the ground state energy.
@@
 From this minimization we can obtain, following standard orbital rotation techniques [70, 62, 68], a new Fock operator with matrix elements:
-where are either positron or electron indices. Hence the Fock operator in block-matrix form is the following:
-where indices () correspond to occupied(unoccupied) electronic spinors.
-Barred indices () refer instead to positronics orbitals.
-The Fock operator in Eq.
-68 can be applied in a Roothan-Hall like procedure [63]:
+where are either positron or electron indices. The Fock operator in Eq. 68 can be applied in a Roothan-Hall like procedure [70]:
 to optimize the orbital coefficients.
@@
 In standard relativistic quantum chemistry, the positrons are usually removed from the treatment by application of the no-pair approximation [62, 53].
-Consequently, the spinor optimization procedure is in fact an excited state problem where the ground state energy corresponds to the first positive energy state.
-This is referred to as a minimax problem [57, 45].
+Consequently, the spinor optimization procedure is in fact a minimax problem [62, 53].
 In our approach, equivalent results can be obtained when no explicit positron is included in the wave function ansatz:
@@
 II.3.2 Kinetic Balance in presence of the field
-In relativistic quantum chemistry, without including positrons, the spinors can be expressed as two-components objects:
+In relativistic quantum chemistry, the spinors can be expressed as two-components objects:
 where and are called respectively ”Large” and ”Small” components.
 In practice, the spinor solutions are expanded on a basis-set: where and is a 2-spinor.
-In order to avoid the so-called varitional collapse of the solution, a constraint on the large and small components of the basis set is applied [53, 45].
+In order to avoid the so-called variational collapse of the solution, a constraint on the large and small components of the basis set is applied [58, 53].
 In the presence of a vector potential, such constraint has the following form:
@@
 We refer to this condition as the kinetic balance [58, 53].
-In our case, since the generalized momentum is a priori depending on the field, the magnetic balance needs to be in principle satisfied.
-This could require to have a field-dependent small-component.
-However, in the present context, we have shown that in the dipole-approximation, applying the length gauge transformation, the momentum of the theory is therefore transformed in the following way (see App.A):
+In our case, since the generalized momentum is a priori depending on the field, the magnetic balance needs, in principle, to be satisfied.
+This could require to have a field-dependent small-component basis-set.
+However, in the present context, we have shown that in the dipole-approximation, applying the length gauge transformation, the momentum of the theory is thus transformed in the following way (see App.A):
 Consequently, the kinetic balance condition:
 can be applied in this case.
-It is important to stress that this is only possible under the dipole-approximation, otherwise, the length gauge transformation would not provide such a simple shape for the Hamiltonian and the associated momentum.
+It is important to stress that this is only possible under the dipole-approximation, otherwise, the length gauge transformation would not provide such a simple expression for the Hamiltonian and the associated momentum.
 II.4 Excited states properties
 In polaritonic chemistry, the signature property emerging from the strong coupling condition is the Rabi-splitting.
-It represents the energy separations between the polaritons formed by the mixing between matter and field states (see Fig.
+It represents the energy separation between the polaritons formed by the mixing between matter and field states (see Fig.
 2).
-Calculating Rabi splittings, which could directly be compared with experimental data, requires access to the excited states of the coupled light matter system.
+Calculating Rabi-splittings, which could directly be compared with experimental data, requires access to the excited states of the coupled light matter system.
 At the HF level they can be simulated recurring to linear response theory.
@@
 [75] presented an HF linear response theory for non-relativistic polaritonic systems.
-In this section, we present an extension of this approach to 4-components Dirac HF.Using linear response theory, excitations energies () can be obtained by solving the well known Casida equation [69]:
-The polaritonic espressions of the and matrices as derived in Ref. [68] assumes the form:
+In this section, we present an extension of this approach to 4-components Dirac HF.
+Using linear response theory, excitations energies () can be obtained by solving the well known Casida equation [76]:
+The polaritonic expressions of the and matrices as derived in Ref. [75] assumes the form:
 where and have matrix elements:
-A similar approach, has been recently proposed by Konecny et al. [30] to calculate excitation energies at the QED Dirac Kohn Sham level of theory.
+A similar approach, has been recently proposed by Konecny et al.
+[35] to calculate excitation energies at the QED Dirac Kohn Sham level of theory based on previous work presented in the following references [77, 78].
 III Results
@@
 The coupling value has been set to 0.05 a.u..
+The datasets generated and analyzed during the current study can be reproduced using the PySCF development code, the inputs files and the geometries provided at the Zenodo link [86].
 III.1 Ground state properties
@@
 In fact, gold complexes are well known to have peculiar electronic properties due to the significant relativistic effects [87, 88, 89, 90].
-In this paper, we will analyze how the polaritonic effects compete with the relativistic ones generating modification of the molecular electronic structure.
+In this paper, we will analyze how the polaritonic effects compete with the relativistic ones, generating modification of the molecular electronic structure.
 In Table 1 we compare the total energies calculated at the DHF level of theory with those evaluated including the field (Pol-DHF).
@@
 | AuH | -21639.06979 | -21639.07709 | -0.1988 |
-The field induced effects, despite small compared to the total energies, represent still a sizable (some tens of eV) variation on the energy of the system.
+The field induced effects, despite small compared to the total energies, still represent a sizable (some tens of eV) variation on the energy of the system.
 In particular, it is interesting noticing that the field induced energy variation is quite similar for copper and gold (Cu - 0.15 eV, Au - 0.20 eV) while it is slightly bigger (in absolute value) for silver (Ag - 0.24 eV).
-This trend could be actually explained if we realize that under the QED-HF approximation, the only field term contributing to the total energy of the system is the dipole self-energy contribution .
+This trend can actually be explained by the fact that under the QED-HF approximation, the only field term contributing to the total energy of the system is the dipole self-energy contribution .
 As highlighted in Eq.
@@
 The effects observed on the total energy values are an indirect observation of the variations induced by the field on the system’s molecular orbitals (MO).
-In Figure 3 we compare the orbital energies for the valence orbitals ( and ) of the three complexes evaluated at the Pol-DHF level with those calculated at the DHF and scalar 2-components X2C level of theory [83, 70, 71].
+In Figure 3, we compare the orbital energies for the valence orbitals ( and ) of the three complexes evaluated at the Pol-DHF level with those calculated at the DHF and scalar 2-components X2C level of theory [93, 79, 80].
 As expected, the biggest variations in the orbital energies can be observed comparing the scalar X2C results with the DHF ones.
@@
 The reduction of the gap between these occupied MOs decreases with the increasing relativistic character of the metal (smaller for Au than Cu).
-Much smaller, even if sizable, is the effect produced by the field.
+The effect of the field on the electronic structure, though sizeable, is much smaller.
 | molecule | (meV) | (%) |
@@
 | AuH | 11.2 | 0.98 |
-The field induce effects () on the energy gap, barely appreciable in Figure 3, are reported in Table 2.
-As already observed for the total energy also in the field induced variation of the energy gap is larger for AgH compared to the CuH and, in particular, to AuH.
-It is crucial, to point out that also in the gold case (smallest effect) the cavity field is able to induce a 10 meV variation on the orbital gap.
+The field induced effects () on the energy gap, barely appreciable in Figure 3, are reported in Table 2.
+As already observed for the total energy, the largest field induced variation of the energy gap is observed for AgH.
+It is crucial, to point out that even in the gold case, where the effect is smaller, the cavity field is able to induce a 10 meV variation on the orbital gap.
 These values, represent a reasonable fraction ( or higher) of a kcal/mol, relevant to observe variations of the chemical properties.
 Even more interesting is observing that the relative change induced by the field () is larger for AuH (about 1% of the total gap) compared to the other systems (0.7% for AgH and 0.6% for CuH).
-Notice that these field induced variations of the energy gap bring (for all the systems) to a reduction of the gap due to a very small stabilization of the MOs accompanied by a larger destabilization of the orbitals.
+Notice that these field induced variations of the energy gap lead (for all systems) to a reduction of the gap due to a very small stabilization of the MOs accompanied by a larger destabilization of the orbitals.
 III.1.1 Gaunt, Breit and polaritonic contributions to ground state energies
-As mentioned in section II.2.1 the radiative QED corrections and in particular the retardation effects can compete or sum up to the polaritonic contributions in a relativistic molecular system.
+As mentioned in section II.2.1, the retardation effects can compete or sum up to the polaritonic contributions in a relativistic molecular system.
 Here we perform a detailed comparison between the energy variations introduced by quantum field, with the Gaunt and Breit terms contributing to the full retarded potential.
@@
 | AuH | 544.0871 | -66.5148 | 544.0874 | -66.5149 |
-In Table 3 the Gaunt/Breit corrections to the DHF and Pol-DHF energies are presented.
-These effects are quite sizable, in particular for AuH that is the system showing the largest relativistic effects.
-In this case the trend is monotonic and both the Gaunt and Breit effect increase going down the group of the periodic table.
+In Table 3, the Gaunt/Breit corrections to the DHF and Pol-DHF energies are presented.
+These effects are quite sizable, in particular for AuH which is the system exhibiting the largest relativistic effects.
+In this case, the trend is monotonic and both the Gaunt and Breit effect increase going down the group of the periodic table.
 It is important to notice that if the absolute energy variation is taken into account, these effects are at least one order of magnitude larger than the effects generated by the quantum field.
-Emblematic is the comparison between the Breit correction in CuH (the less relativistic system) and the corresponding polaritonic energy correction, shown in Table 1.
-These observations clearly indicate that, for a molecular system coupled to a quantum field, attention needs to be paid in omitting the retardation effects, a generic pratice in standard electronic relativistic quantum chemistry simultations.More striking results can be obtained if we perform the same analysis on the orbital energies.
-In Figure 4 the energy contributions (in logaritmic scale) due to the Gaunt, Breit and polaritonic terms on the energies of the occupied MOs are presented for the three systems.
+The comparison between the Breit correction in CuH (the less relativistic system) and the corresponding polaritonic energy correction, shown in Table 1, demonstrates that quite unequivocally.
+These observations clearly indicate that, for a molecular system coupled to a quantum field, attention needs to be paid in omitting the retardation effects, a generic practice in standard electronic relativistic quantum chemistry simulations.
+More striking results can be obtained if we perform the same analysis on the orbital energies.
+In Figure 4, the absolute value of the energy contributions (in logarithmic scale) due to the Gaunt, Breit and polaritonic terms on the energies of the occupied MOs are presented for the three systems.
 The most evident aspect is the very wide variation range of the energy contributions for different molecular orbitals.
 In general, the effects due to the full-Breit term are much more sizable for the core orbitals, while they monotonically decrease moving toward the valence ones.
-This trend is clearly expected, since the inner orbitals host the fastest electrons that are more affected by the relativistic effects.
+This trend is clearly expected and has been documented in Ref.
+[63].
 This behavior is obviously more evident for AuH than for CuH.
@@
 In the core region, the retardation effects are always larger, while in the valence region the polaritonic effects are dominant.
-It is interesting highlighting that for the intermediate orbitals all the effects are quite comparable, confirming that the inclusion of retardation effects could produce crucial effects in molecular systems coupled to photons.
-There are anyway cases where, neglecting such effects, as usually done in relativistic quantum chemistry, is still reasonable.
+Interestingly, for the intermediate orbitals, all the effects are quite comparable, confirming that the inclusion of retardation effects could induce crucial effects in molecular systems coupled to photons.
+Still, there are cases where neglecting such effects, as usually done in relativistic quantum chemistry, is still reasonable.
 For instance, if we are interested in excited states properties (i.e.
@@
 in Ref.
-[30], then omitting retardation effects in the treatment is clearly acceptable.
+[35], then omitting retardation effects in the treatment is clearly meaningful.
+For this reason, in the next section excited states will be investigated without including the full-Breit correction.
 However, if excitations from inner orbitals (i.e.
-core excitations, etc.) or magnetic properties of the system need to be analyzed the inclusion of the retardation effects will become a crucial point in the treatment.
-It is important to point out that the implications of the effects presented in this section are, in some sense, minimized by the absence of electron-electron and electron-photon correlation in the treatment.
+core excitations, etc.) or magnetic properties of the system need to be analyzed, the inclusion of the retardation effects will become necessary.
+Bear in mind that the implications of the effects presented in this section are, in some sense, minimized by the absence of electron-electron and electron-photon correlation in the treatment.
 For instance, the HF approximation removes the frequency dependence of the ground state energy from the frequency of the field.
-This dependence can be recovered only by including correlation into the model.
+This dependence can be recovered only by including correlation into the model [27].
 We expect, that the inclusion of frequency dependent terms, will be crucial in particular to describe resonant processes.
@@
 Similar analysis for CuH and AgH is presented in Appendix B.
-In this system, because of strong relativistic effects, we expect to observe spectra very different compared to those that can be simulated without taking relativity into account.
-This is evident if we look at Figure 5.
+In this system, because of strong relativistic effects, we expect to observe spectra very different compared to those simulated without taking relativity into account.
+This is obvious looking at Figure 5.
 As it can be seen, at the HF level, the triplet state () is found to be higher in energy than the singlets states, in disagreement with the experimental data from Ref.s [94, 95].
@@
 This choice highlights the energy shift due to the dipole self energy.
-In resonance conditions also the Rabi splitting (discussed later in Figure 6) would be visible.
-Comparing this data with the bare electronic DHF ones, we see that the field (in off resonance conditions) induces a stabilization of the and states but without inducing any change in the ordering of the states.
+In resonance conditions, the Rabi splitting (discussed later in Figure 6) would also be visible.
+Comparing this data with the bare electronic DHF ones, we see that the field (in off resonance conditions) induces a stabilization of the and states without inducing any change in the ordering of the states.
 Finally, in Fig.6 we have reported the dispersion of the AuH excitation energies as a function of the cavity frequency.
@@
 For this system, the energy difference between and , associated to spin-orbit coupling, is significantly larger (0.04 eV) than the Rabi-splitting.
-This spin-orbit coupling induced singlet-triplet rabi-splitting was already reported for a different system by Koneckny et al.
+This spin-orbit coupling induced singlet-triplet Rabi-splitting was already reported for a different system by Konecny et al.
 in Ref.
 [35].
-These observations clearly prove how the electromagnetic field can be used to manipulate and control inter-system crossing processes and consequently the phosphorescence of complexes containing heavy atoms.
+These observations clearly demonstrate how the electromagnetic field can be used to manipulate and control inter-system crossing processes and consequently the phosphorescence of complexes containing heavy atoms.
 In App.
-B a similar discussion is reported also for CuH and AgH.
+B a similar discussion is also reported for CuH and AgH.
 IV Conclusions
-In this paper we proposed a reformulation of relativistic QED allowing for an easier development of ab initio methodologies to simulate heavy atoms molecular complexes in strong coupling conditions.
+In this paper, we proposed a reformulation of relativistic QED allowing for an easier development of ab initio methodologies to simulate heavy atoms molecular complexes in strong coupling conditions.
 Using this theoretical ground, we reported the development and implementation of the first relativistic polaritonic wave function based ab initio method, namely Pol-DHF.
 The theory has been presented starting from the standard Lagrangian and has been derived into a usable implementation of the Pol-DHF code.
-Considering the possible competition with the polaritonic effects, the radiative QED corrections have been explicitly included into the treatment.
-After providing our roadmap to the implementation, we presented applications of Pol-DHF to three metal hydrides: CuH, AgH and, AuH.
+Considering the possible competition with the polaritonic effects, the importance of radiative QED corrections have been addressed.
+After providing our road-map to the implementation, we presented applications of Pol-DHF to three metal hydrides: CuH, AgH and, AuH.
 These systems were an excellent test case in order to assess the magnitude of the polaritonic effects in comparison to the relativistic effects.
@@
 The polaritonic effects resulted having the largest relative influence on AuH for which relativistic effects are more prominent.
-Afterwards, we provided a detailed analysis of the competition between the Gaunt, Breit and polaritonic effects on the ground state and orbital energies.
+Afterward, we provided a detailed analysis of the competition between the Gaunt, Breit and polaritonic effects on the ground state and orbital energies.
 We have shown that even though the full-Breit term represents a significantly larger contribution to the total ground state energy, its effect is extremely orbital dependent.
-In particular, it is much larger on the core orbitals compare to the valence, where instead the polaritonic effects dominate.
-Therefore, we could conclude that neglecting the Full-Breit term (practice usually applied in relativistic quantum chemistry) in strong coupling conditions is only possible if properties involving valence orbitals need to be investigated.
+In particular, its impact is much stronger on the core orbitals compared to the valence, where instead the polaritonic effects dominate.
+Therefore, we could conclude that neglecting the full-Breit term (a common practice in relativistic quantum chemistry) in strong coupling conditions is only possible if properties involving exclusively valence orbitals need to be investigated.
+A similar study could be conducted adding effective potentials to include radiative QED corrections and comparing the magnitude of their effects to the polaritonic ones, this will surely be the topic of a future work.
 Lastly, we presented excited state calculations for AuH at the TD-Pol-DHF level.
-As already observed by Koneckny et al.
-[30], we have shown that using a fully relativistic polaritonic theory the appearance of Rabi-Splittings at the crossing between singlet and triplet potential energy surfaces can be observed.
+As already observed by Konecny et al.
+[35], we have shown that using a fully relativistic polaritonic theory, the appearance of Rabi-Splittings at the crossing between singlet and triplet potential energy surfaces can be observed.
 These results clearly open new possibilities for the control of the inter-system crossing in photochemistry.
-We strongly believe that the methodologies and the applications presented in this paper will represent a groundbreaking step toward the simulation of relativistic molecular systems strongly coupled to photons, field that is recently finding many interesting applications in photochemistry but also in spintronics and quantum computation.
+We strongly believe that the methodologies and the applications presented in this paper can represent a groundbreaking step toward the simulation of relativistic molecular systems strongly coupled to photons, field that is recently finding many interesting applications in photochemistry but also in spintronics and quantum computation.
 Acknowledgements.
-Guillaume Thiam, Riccardo Rossi and Enrico Ronca acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Program (Grant ERC-StG-2021-101040197-QED-SPIN).
-Henrik Koch acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (Grant Agreement 101020016).Appendix A Length gauge transformation
+The authors acknowledge Michael Ruggenthaler, Lukas Konecny, Rosario Roberto Riso and Giovanni Bistoni for insightful discussions.
+G.
+T., R.
+R.
+and E.
+R.
+acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Program (Grant ERC-StG-2021-101040197-QED-SPIN).
+H.
+K.
+acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (Grant Agreement 101020016).Appendix A Length gauge transformation
 In section II.2.3, we present the results of the application of the length gauge transformation.
@@
 It is important to emphasize that the dipole approximation strongly simplifies the expression of the terms involving the momentum.
-The other noticeable fact is that the terms involving cancels out (the translated momentum bring out a term cancelling out with the unmodified term).
+The other noticeable fact is that the terms involving cancels out (the translated momentum bring out a term canceling out with the unmodified term).
 Let us now look at the modification of the photon number operator. The Becker-Campbell-Hausdorff formula implies that: